OpenStudy (anonymous):

what is the value of y so that the ordered pair (5, y) is a solution to the equation -5x + y = -40

OpenStudy (abdullahm):

Plug in x = 5, and solve for y.

OpenStudy (abdullahm):

-5 (5) + y = 40 y = ??

OpenStudy (abdullahm):

-5 (5) + y = 40 y = ??

OpenStudy (anonymous):

35

OpenStudy (abdullahm):

What is -5 multiplied by 5?

OpenStudy (anonymous):

-25

OpenStudy (abdullahm):

Hint: -5 * 5 = (-1) * (5 * 5) = ??

OpenStudy (abdullahm):

Ok, so now we have: -25 + y = 40 We need to isolate y. So if we add 25 on both sides we will get y all alone. So, add 25 on both sides. What do you get?

OpenStudy (anonymous):

25 and 65

OpenStudy (abdullahm):

My bad, I keep forgetting that it's -40 -25 + y = -40 y = -40 + (-25) = ??

OpenStudy (anonymous):

-65

OpenStudy (abdullahm):

My bad, I keep forgetting that it's -40 -25 + y = -40 y = -40 + (-25) = ??

OpenStudy (anonymous):

-65

OpenStudy (abdullahm):

Errr, I keep making a mistake. y = -40 + 25 = ??

OpenStudy (abdullahm):

This will get you your final correct answer. :)

OpenStudy (mathmale):

You are saying that y=-65. Check your own work by substituting x=5 and y=-65 into the given equation. Either this equation will be true or it will be false. Which result do you get?

OpenStudy (anonymous):

-25

OpenStudy (abdullahm):

Are you sure? -50 + 25 = -25 What is -40 + 25 = ?

OpenStudy (anonymous):

-15

OpenStudy (abdullahm):

And that's your final answer!

OpenStudy (anonymous):

thank you

OpenStudy (abdullahm):

It was my pleasure!